A ball thrown by one player to another is in air for 2 s. The maximum height attained by the ball is (g=10 m/s2)

Grade:12th pass

1 Answers

Nehal
6 years ago
Let u be the initial velocity of the balllUsing equation of a projectile motionTime period in which the body remains in air isT=2u sin90/g=2u=10Maximum height attained by the body is H=u^2/2gH=5